An angry construction worker throws his wrench downward from a height of 128 feet
with an initial velocity of 32 feet per second. The height of the wrench above the ground after t seconds is given by S(t)= - 16t - 32t - 128.
a) What is the height of the wrench after 1 second?
b) How long does it take for the wrench to reach the
ground?

Answer by ankor@dixie-net.com(22740)

wrench thrown downward from a height of 128 feet with an initial velocity
of 32 feet per second. The height of the wrench above the ground after
t seconds is given by S(t)= -16t^2 - 32t - 128.
The formula should be -16t^2 (gravity)
;
a) What is the height of the wrench after 1 second?
Substitute 1 for t and find S(t)
S(t) = -16(1^2) - 32(1) + 128
s(t) = -16 - 32 + 128
S(t) = 80 ft is the height of the wrench after 1 second
;

b) How long does it take for the wrench to reach the ground?
:
The ground is S(t) = 0, we want to find t, so we have:
-16t^2 - 32t + 128 = 0
Simplify, divide equation by -16, this also changes the signs
t^2 + 2t - 8 = 0
Factor this to
(t+4)(t-2) = 0
The positive solution:
t = 2 seconds for the wrench to reach the ground